# Geometric structures on the quaternionic unit ball and slice regular Möbius transformations

### v1

@@ Abstract.
Building from ideas of hypercomplex analysis on the quaternionic unit ball, we introduce Hermitian, Riemannian and Kähler-like structures on the latter. These are built from the so-called regular Möbius transformations. Such geometric structures are shown to be natural generalizations of those from the complex setup. Our structures can be considered as more natural, from the analytic viewpoint, than the usual quaternionic hyperbolic geometry. Furthermore, our constructions provide solutions to problems not achieved by hyper-Kähler and quaternion-Kähler geometries when applied to the quaternionic unit ball.
@@ 1. Introduction
One might expect that the coincidence between the $4$-dimensional real and $1$-dimensional quaternionic hyperbolic geometries can also be related to a suitable function theory supported by the quaternions $\mathbb{H}$. However, it is well known that a naive straightforward generalization of holomorphicity from $\mathbb{C}$ to $\mathbb{H}$ does not yield the desired results, mainly due to the non-commutativity of $\mathbb{H}$. Nevertheless, it has also been shown that the introduction of appropriate notions of regularity or “hyperholomorphicity” produce very useful function theories on $\mathbb{H}$. In this work we will focus on the notion of slice or Cullen regularity deeply studied in [10] (see also [9]). Besides allowing to generalize many of the classical complex theorems, slice regularity has a natural characterization: on unit balls centered at the origin, slice regular functions are precisely those with a power series expansion with right coefficients. Furthermore, the current literature shows that a mature and very well developed so-called hypercomplex analysis is readily available through the use of slice regularity and their extensions. As an example, we refer to [6, 8] for corresponding extensions in the case of Clifford algebras.
@@
On the other hand, the isometries of the $1$-dimensional quaternionic hyperbolic geometry are given by suitable linear fractional transformations of $\mathbb{H}$, which turn out to be non-regular. Nevertheless, using the natural so-called regular product, [13] developed the notion of regular linear fractional transformations, also referred as regular Möbius transformations. These are the natural ones to consider from the viewpoint of hypercomplex analysis. Hence, on the unit ball $\mathbb{B}$ of $\mathbb{H}$ centered at the origin we have two sorts of Möbius transformations, regular and non-regular, while $\mathbb{B}$ carries the metric realizing both the $4$-dimensional real and $1$-dimensional quaternionic hyperbolic geometries. The non-regular Möbius transformations preserve such hyperbolic geometry (see [2, 4, 5]) but does not relate properly to hypercomplex analysis. The regular Möbius transformations do relate nicely to the latter but, as shown in [3], regular Möbius transformations do not preserve the hyperbolic geometry. We thus arrive to the seemingly impossibility of studying a hypercomplex hyperbolic function theory. However, [3] proposed the problem of a further study of regular Möbius transformations and the geometry of $\mathbb{B}$. As a matter of fact, it would be interesting to study as well the Clifford algebra case considered in [7]. However, our focus will lie in the quaternionic unit ball.
@@
The goal of this work is to present a solution to the problems formulated so far on the analysis and geometry of $\mathbb{B}$. We introduce in section 5 what we call slice geometric structures on $\mathbb{B}$, which include Hermitian, Riemannian and Kähler geometries of a truly quaternionic nature (see Definition 5.6). These are constructed out of regular Möbius transformations in a similar way as the corresponding structures on the unit disk on the complex plane are obtained from (complex holomorphic) Möbius transformations. Our slice geometric structures are computed explicitly and some fundamental properties are obtained in section 6, see Theorems 5.7, 6.1 and 6.4, thus verifying that they satisfy their claimed nature.
@@
On the other hand, an important feature of the hypercomplex analysis of slice regular functions are the so-called representation formulas: every regular function can be determined from its values on a single slice. We prove in Corollaries 6.3 and 6.7 that our three slice geometric structures, Riemannian, Hermitian and Kähler, satisfy similar representation formulas, i.e. they can be recovered from a single slice. Furthermore, our formulas can be seen as invariant, in the Riemannian case, and equivariant, in the Hermitian and Kähler cases, identities with respect to the action by conjugation of the group of unit quaternions. The latter is by itself an interesting set of facts from the viewpoint of Lie theory.
We recall that differential geometry has already provided an approach to the quaternionic setup: hyper-Kähler and quaternion-Kähler manifolds. We discuss in Remark 6.8 the relationship between those notions and our slice regular approach. As noted there, our methods are better fit to both the geometry and hypercomplex analysis on $\mathbb{B}$. In particular, we observe that $\mathbb{B}$, with the usual hyperbolic metric $\widehat{G}$, is not hyper-Kähler and, although it is quaternion-Kähler, it does not provide globally defined $2$-forms. Our approach does yield both a Kähler-like structure and $2$-forms associated to a Riemannian metric $G$ nicely adapted to hypercomplex analysis.
As for the distribution of this work, sections 2 and 3 introduce our two main tools: slice regularity and Möbius transformations, both non-regular and regular. The brief section 4 recalls the construction of the geometric structures on the complex unit disk. As described above, sections 5 and 6 contains the main constructions and results.
The author would like to thank Cinzia Bisi and Caterina Stoppato for their input, which allowed to improve this work.
@@ 2. Slice regular functions
We will denote from now on by $\mathbb{H}$ the division real algebra of quaternions and by $\mathbb{B}$ its unit ball centered at the origin. As usual $\operatorname{Re}$ and $\operatorname{Im}$ will denote the real part and imaginary part functions defined on $\mathbb{H}$. In particular, $\operatorname{Im}(\mathbb{H})$ denotes the space of purely imaginary quaternions so that its subset $\mathbb{S}$, consisting of the unitary purely imaginary quaternions, is the $2$-dimensional sphere of imaginary units of $\mathbb{H}$. For every $I\in\mathbb{S}$, we will denote by $\mathbb{C}_{I}$ the complex plane in $\mathbb{H}$ whose elements are of the form $x+yI$, where $x,y\in\mathbb{R}$. Correspondingly, the slice of $\mathbb{B}$ determined by a given $I\in\mathbb{S}$ will be denoted by $\mathbb{D}_{I}=\mathbb{B}\cap\mathbb{C}_{I}$, which is the unit disk in $\mathbb{C}_{I}$. We will also consider the complex plane $\mathbb{C}$ in the classical sense and denote by $\mathbb{D}$ its unit disk centered at the origin. Note that, in this work and for the case of $\mathbb{C}$ (not considered as contained in $\mathbb{H}$), the real and imaginary parts will have their classical meaning as real numbers. This is important to keep in mind since we will have the occasion to deal with some classical geometric objects.
The fundamental notion of slice regularity as first formulated in [10] (see also [9]) is introduced in the following definition.
@@
A function $f\mathrel{\mathop{\ordinarycolon}}\mathbb{B}\rightarrow\mathbb{H}$ is called slice regular if the equation
@@
The theory of slice regular functions on $\mathbb{B}$, and furthermore on suitable domains of $\mathbb{H}$, has been deeply developed. We will use well known facts from this theory and refer to [9] for further details. However, we will have the occasion to provide a precise references when this helps to clarify our discussion.
Every slice regular function $f\mathrel{\mathop{\ordinarycolon}}\mathbb{B}\rightarrow\mathbb{H}$ admits a power series expansion
@@
with coefficients in $\mathbb{H}$ that converges uniformly on compact sets. This is used to define a product of slice regular functions on $\mathbb{B}$.
@@
For $f,g\mathrel{\mathop{\ordinarycolon}}\mathbb{B}\rightarrow\mathbb{H}$ slice regular functions admitting power series expansions
@@
the (slice regular) ***-product is the function $f*g\mathrel{\mathop{\ordinarycolon}}\mathbb{B}\rightarrow\mathbb{H}$ given by
@@
It is a well known fact that the ***-product of slice regular functions is slice regular, and also that the space of all such functions becomes a real algebra with the ***-product. Furthermore, one can define reciprocals with respect to the ***-product as follows.
@@
Let $f\mathrel{\mathop{\ordinarycolon}}\mathbb{B}\rightarrow\mathbb{H}$ be a slice regular function with power series expansion given by
@@
The regular conjugate $f^{c}\mathrel{\mathop{\ordinarycolon}}\mathbb{B}\rightarrow\mathbb{H}$ is the (slice regular) function given by
$f(q)=\sum_{n=0}^{\infty}q^{n}\overline{a}_{n}.$
The symmetrization $f^{s}$ of $f$ is the (slice regular) function defined as $f^{s}=f^{c}*f=f*f^{c}$. The regular reciprocal of $f$ is the function $f^{-*}\mathrel{\mathop{\ordinarycolon}}\mathbb{B}\setminus Z_{f^{s}}\rightarrow\mathbb{H}$ defined as
@@ 3. Linear fractional transformations
the matrix that yields pseudo-Hermitian inner products of signature $(1,1)$ on both $\mathbb{C}^{2}$ and $\mathbb{H}^{2}$. The corresponding isometry groups for such inner products are
$\mathrm{U}(1,1)$ =$=\{A\in M_{2}(\mathbb{C})\mid AI_{1,1}A^{*}=I_{1,1}\},$
$\mathrm{Sp}(1,1)$ =$=\{A\in M_{2}(\mathbb{H})\mid AI_{1,1}A^{*}=I_{1,1}\},$
called the pseudo-unitary, complex and quaternionic, respectively, Lie groups of signature $(1,1)$. We also have the Lie groups $\mathbb{T}$ and $\mathrm{Sp}(1)$ which consist of the complex and quaternionic elements, respectively, with norm $1$. In particular, both are the unit spheres centered at the origin in their corresponding division real algebras. Finally we will denote by $\mathbb{T}\times\mathbb{T}$ and $\mathrm{Sp}(1)\times\mathrm{Sp}(1)$ the groups of $2\times 2$ diagonal matrices with entries in $\mathbb{T}$ and $\mathrm{Sp}(1)$ in the diagonal, respectively. It is straightforward to check that $\mathbb{T}\times\mathbb{T}\subset\mathrm{U}(1,1)$ and $\mathrm{Sp}(1)\times\mathrm{Sp}(1)\subset\mathrm{Sp}(1,1)$. Moreover, both are maximal compact subgroups (see [12]).
@@
It is a known fact that the action of $\mathrm{Sp}(1,1)$ on $\mathbb{B}$ defined above realizes all linear fractional transformations, with right coefficients, that map $\mathbb{B}$ onto itself. We refer to [2, 4, 5] for the proof of this claim. We observe that in most of the previous works dealing with quaternionic Möbius transformations, with the exception of [2, 3], one considers coefficients on the left which lead to a corresponding left action. We have chosen to use coefficients and corresponding action on the right to better resemble slice regular maps. Left multiplication by a constant quaternion is not slice regular but right multiplication is.
On the other hand, [13] introduced the regular counterpart of linear fractional transformations (see also [9]). For every $A\in\mathrm{Sp}(1,1)$ as above the regular Möbius transformation of $\mathbb{B}$ associated to $A$ is the (regular) map $\mathcal{F}_{A}\mathrel{\mathop{\ordinarycolon}}\mathbb{B}\rightarrow\mathbb{B}$ given by
@@
Every regular Möbius transformation is a homeomorphism of $\mathbb{B}$. However, as a consequence of the properties of the regular product, the assignment $A\mapsto\mathcal{F}_{A}$ does not define an action. Nevertheless, these are the transformations that better fit the analysis of slice regular functions. We will denote by $\mathcal{M}(\mathbb{B})$ the set of all regular Möbius transformations. In particular, $\mathcal{M}(\mathbb{B})$ is not a group for the usual composition of functions. Whenever we need to clearly distinguish the transformations of $\mathcal{M}(\mathcal{B})$ from those obtained from the $\mathrm{Sp}(1,1)$-action considered above, we will refer to the latter as non-regular Möbius transformations.
@@
A map :$\mathcal{F}\mathrel{\mathop{\ordinarycolon}}\mathbb{B}\rightarrow\mathbb{B}$ is a regular Möbius transformation if and only if there exist $a\in\mathbb{B}$ and $u\in\mathrm{Sp}(1)$ such that
@@
    $\mathcal{F}_{a}$ is a diffeomorphism whose differential at $a$ is given by
    $\operatorname{d\!}\>(\mathcal{F}_{a})_{a}(\alpha)=\frac{(1-a^{2})^{-1}(a\alpha a-\alpha)}{1-|a|^{2}},$
@@
By definition we have
$(q\overline{a}-1)^{s}$ $=(q\overline{a}-1)*(qa-1)=q^{2}|a|^{2}-2q\operatorname{Re}(a)+1$
$(qa-1)*(q-a)$ $=q^{2}a-q(a^{2}+1)+a$
and so the expression for $\mathcal{F}_{a}$ is now immediate. The statement (1) now follows from direct substitution. The first claim in (2) is a consequence of Corollary 8.24 and Theorem 9.12 from [9]. By the former $\mathcal{F}_{a}$ is a local diffeomorphism and by the latter $\mathcal{F}_{a}$ is a homeomorphism. Finally, the formula in (2) follows from the direct computation of the derivative at $t=0$ of the map $t\mapsto\mathcal{F}_{a}(a+t\alpha)$. ∎

### v2

@@ Abstract.
Building from ideas of hypercomplex analysis on the quaternionic unit ball, we introduce Hermitian, Riemannian and Kähler-like structures on the latter. These are built from the so-called regular Möbius transformations. Such geometric structures are shown to be natural generalizations of those from the complex setup. Our structures can be considered as more natural, from the hypercomplex viewpoint, than the usual quaternionic hyperbolic geometry. Furthermore, our constructions provide solutions to problems not achieved by hyper-Kähler and quaternion-Kähler geometries when applied to the quaternionic unit ball. We prove that the Riemannian metric obtained in this work yields the same tensor previously computed by Arcozzi-Sarfatti. However, our approach is completely geometric as opposed to the function theoretic methods of Arcozzi-Sarfatti.
@@ 1. Introduction
One might expect that the coincidence between the $4$-dimensional real and $1$-dimensional quaternionic hyperbolic geometries can also be related to a suitable function theory supported by the quaternions $\mathbb{H}$. However, it is well known that a naive straightforward generalization of holomorphicity from $\mathbb{C}$ to $\mathbb{H}$ does not yield the desired results, mainly due to the non-commutativity of $\mathbb{H}$. Nevertheless, it has also been shown that the introduction of appropriate notions of regularity or “hyperholomorphicity” produce very useful function theories on $\mathbb{H}$. In this work we will focus on the notion of slice or Cullen regularity deeply studied in [13] (see also [11, 12]). Besides allowing to generalize many of the classical complex theorems, slice regularity has a natural characterization: on unit balls centered at the origin, slice regular functions are precisely those with a power series expansion with right coefficients. Furthermore, the current literature shows that a mature and very well developed so-called hypercomplex analysis is readily available through the use of slice regularity and their extensions. As an example, we refer to [9, 11] for corresponding extensions in the case of Clifford algebras. It is also important to mention [5], where results for self-maps of the complex unit disc are extended to the quaternionic unit ball. This kind of developments are at the heart of this work.
On the other hand, the isometries of the $1$-dimensional quaternionic hyperbolic geometry are given by suitable linear fractional transformations of $\mathbb{H}$, which turn out to be, in general, non-regular. Related to this fact, it was proved in [4] that the geometry on the quaternionic unit ball is not isometric to the Kobayashi geometry on the complex $2$-dimensional unit ball in $\mathbb{C}^{2}$. In other words, there is a marked incompatibility between the quaternionic hyperbolic geometry with $2$-dimensional complex geometries as well.
@@
Nevertheless, using the natural so-called regular product, [16] developed the notion of regular linear fractional transformations, also referred as regular Möbius transformations. These are the natural ones to consider from the viewpoint of hypercomplex analysis. Hence, on the unit ball $\mathbb{B}$ of $\mathbb{H}$ centered at the origin we have two sorts of Möbius transformations, regular and non-regular, while $\mathbb{B}$ carries the metric realizing both the $4$-dimensional real and $1$-dimensional quaternionic hyperbolic geometries. The non-regular Möbius transformations preserve such hyperbolic geometry (see [3, 7, 8]) but does not relate properly to hypercomplex analysis. The regular Möbius transformations do relate nicely to the latter but, as shown in [6], regular Möbius transformations do not preserve the hyperbolic geometry. We thus arrive to the seemingly impossibility of studying a hypercomplex hyperbolic function theory. However, [6] proposed the problem of a further study of regular Möbius transformations and the geometry of $\mathbb{B}$. As a matter of fact, it would be interesting to study as well the Clifford algebra case considered in [10]. However, our focus will lie in the quaternionic unit ball.
We note that [1] has already introduced a Riemannian metric that relates to regular functions. The corresponding geometry was constructed by considering a quaternionic Hardy space, its corresponding reproducing kernel and a pseudo-hyperbolic distance; the latter mirrors the complex case. Finally, it is proved in [1] that such distance yields indeed a Riemannian metric. Furthermore, several non-trivial and interesting geometric properties for this geometry are obtained. On the other hand, the problem of building symplectic and/or Hermitian geometries for $\mathbb{B}$ are not considered in [1].
@@
The goal of this work is to present a solution to the problems formulated so far on the analysis and geometry of $\mathbb{B}$. We introduce in section 5 what we call slice geometric structures on $\mathbb{B}$, which include Hermitian, Riemannian and Kähler-like geometries of a truly quaternionic nature (see Definition 5.6). These are constructed out of regular Möbius transformations in a similar way as the corresponding structures on the unit disk on the complex plane are obtained from (complex holomorphic) Möbius transformations. Our slice geometric structures are computed explicitly and some fundamental properties are obtained in section 6, see Theorems 5.7, 6.1 and 6.7, thus verifying that they satisfy their claimed nature.
@@
Another important feature appears in our development: the Riemannian metric from [1] and our slice Riemannian metric turn out to be the same exact tensor, as it is proved in Theorem 6.4. As noted above, the results from [1] make strong use of function space theoretic methods (see section 6 for some details). Our approach is almost exclusively based on transformations that mirror the group theoretic methods of the complex case. For the quaternionic unit ball $\mathbb{B}$ the regular Möbius transformations is not a group, but we bypass the difficulties involved by taking advantage of the known fundamental properties of such transformations (see [11, 12]).
The fact that the Riemannian metrics constructed in [1] and in this work are the same reveals a fundamental property: both the function space theoretic and the geometric/group-like methods yield the same Riemannian geometry on $\mathbb{B}$. One can argue in favor of the latter method since it turns out to be more straightforward computation-wise. Another contribution of this work is the construction of a Kähler-like structure on $\mathbb{B}$ nicely related to the Riemannian metric that we have discussed.
On the other hand, an important feature of hypercomplex analysis of slice regular functions are the so-called representation formulas: every regular function can be determined from its values on a single slice. We prove in Corollaries 6.6 and 6.10 that our three slice geometric structures, Riemannian, Hermitian and Kähler, satisfy similar representation formulas, i.e. they can be recovered from a single slice. Furthermore, our formulas can be seen as invariant, in the Riemannian case, and equivariant, in the Hermitian and Kähler cases, identities with respect to the action by conjugation of the group of unit quaternions. The latter is by itself an interesting set of facts from the viewpoint of Lie theory.
We recall that differential geometry has already provided an approach to the quaternionic setup: hyper-Kähler and quaternion-Kähler manifolds. We discuss in Remark 6.11 the relationship between those notions and our slice regular approach. As noted there, our methods are better fit to both geometry and hypercomplex analysis on $\mathbb{B}$. In particular, we observe that $\mathbb{B}$, with the usual hyperbolic metric $\widehat{G}$, is not hyper-Kähler and, although it is quaternion-Kähler, it does not provide globally defined $2$-forms. Our approach does yield both a Kähler-like structure and $2$-forms associated to a Riemannian metric $G$ nicely adapted to hypercomplex analysis.
As for the distribution of this work, sections 2 and 3 introduce our two main tools: slice regularity and Möbius transformations, both non-regular and regular. The brief section 4 recalls the construction of the geometric structures on the complex unit disk. As described above, sections 5 and 6 contain the main constructions and results.
The author would like to thank Cinzia Bisi and Caterina Stoppato for their input, which allowed to improve this work. The input from the anonymous reviewers have also provided very useful information that resulted on a much better version of this work.
@@ 2. Slice regular functions
We will denote from now on by $\mathbb{H}$ the division real algebra of quaternions and by $\mathbb{B}$ its unit ball centered at the origin. As usual $\operatorname{Re}$ and $\operatorname{Im}$ will denote the real part and imaginary part functions defined on $\mathbb{H}$. In particular, $\operatorname{Im}(\mathbb{H})$ denotes the space of purely imaginary quaternions so that its subset $\mathbb{S}$, consisting of the unitary purely imaginary quaternions, is the $2$-dimensional sphere of imaginary units of $\mathbb{H}$. For every $I\in\mathbb{S}$, we will denote by $\mathbb{C}_{I}$ the complex plane in $\mathbb{H}$ whose elements are of the form $x+yI$, where $x,y\in\mathbb{R}$. Correspondingly, the slice of $\mathbb{B}$ determined by a given $I\in\mathbb{S}$ will be denoted by $\mathbb{D}_{I}=\mathbb{B}\cap\mathbb{C}_{I}$, which is the unit disk in $\mathbb{C}_{I}$. We will also consider the complex plane $\mathbb{C}$ in the classical sense and denote by $\mathbb{D}$ its unit disk centered at the origin. Note that, in this work and for the case of $\mathbb{C}$ (not considered as contained in $\mathbb{H}$), the real and imaginary parts will have their classical meaning as real numbers. This is important to keep in mind since we will have the occasion to deal with some classical complex geometric objects.
The fundamental notion of slice regularity as first formulated in [13] (see also [11, 12]) is introduced in the following definition.
@@
A function $f:\mathbb{B}\rightarrow\mathbb{H}$ is called slice regular if the equation
@@
The theory of slice regular functions on $\mathbb{B}$, and furthermore on suitable domains of $\mathbb{H}$, has been deeply developed. We will use well known facts from this theory and refer to [12] for further details. However, we will have the occasion to provide precise references when this helps to clarify our discussion.
Every slice regular function $f:\mathbb{B}\rightarrow\mathbb{H}$ admits a power series expansion
@@
with coefficients in $\mathbb{H}$ that converges uniformly on compact sets. Conversely, every such convergent power series yields a slice regular function on $\mathbb{B}$ (see [12, section 1.1]). This allows us to introduce the following constructions for slice regular functions on $\mathbb{B}$. We refer to [11, 12] for further details and the proofs of our claims below.
@@
For $f,g:\mathbb{B}\rightarrow\mathbb{H}$ slice regular functions admitting power series expansions
@@
the (slice regular) *-product is the function $f*g:\mathbb{B}\rightarrow\mathbb{H}$ given by
@@
It is a well known fact that the *-product of slice regular functions is slice regular, and also that the space of all such functions becomes a real algebra with the *-product. Furthermore, one can define reciprocals with respect to the *-product as follows.
@@
Let $f:\mathbb{B}\rightarrow\mathbb{H}$ be a slice regular function with power series expansion given by
@@
The regular conjugate $f^{c}:\mathbb{B}\rightarrow\mathbb{H}$ is the (slice regular) function given by
$f^{c}(q)=\sum_{n=0}^{\infty}q^{n}\overline{a}_{n}.$
The symmetrization $f^{s}$ of $f$ is the (slice regular) function defined as $f^{s}=f^{c}*f=f*f^{c}$. The regular reciprocal of $f$ is the function $f^{-*}:\mathbb{B}\setminus Z_{f^{s}}\rightarrow\mathbb{H}$ defined as
@@ 3. Linear fractional transformations
the matrix that yields pseudo-Hermitian inner products of signature $(1,1)$ on both $\mathbb{C}^{2}$ and $\mathbb{H}^{2}$. The corresponding isometry groups for those inner products are
$\mathrm{U}(1,1)$ =$=\{A\in M_{2}(\mathbb{C})\mid A^{*}I_{1,1}A=I_{1,1}\},$
$\mathrm{Sp}(1,1)$ =$=\{A\in M_{2}(\mathbb{H})\mid A^{*}I_{1,1}A=I_{1,1}\},$
called the pseudo-unitary, complex and quaternionic, respectively, Lie groups of signature $(1,1)$. We also have the Lie groups $\mathbb{T}$ and $\mathrm{Sp}(1)$ which consist of the complex and quaternionic elements, respectively, with norm $1$. In particular, both are the unit spheres centered at the origin in their corresponding division real algebras. Finally, we will denote by $\mathbb{T}\times\mathbb{T}$ and $\mathrm{Sp}(1)\times\mathrm{Sp}(1)$ the groups of $2\times 2$ diagonal matrices with entries in $\mathbb{T}$ and $\mathrm{Sp}(1)$ in the diagonal, respectively. It is straightforward to check that $\mathbb{T}\times\mathbb{T}\subset\mathrm{U}(1,1)$ and $\mathrm{Sp}(1)\times\mathrm{Sp}(1)\subset\mathrm{Sp}(1,1)$. Moreover, both are maximal compact subgroups (see [15]).
@@
It is a known fact that the action of $\mathrm{Sp}(1,1)$ on $\mathbb{B}$ defined above realizes all linear fractional transformations, with right coefficients, that map $\mathbb{B}$ onto itself. We refer to [3, 7, 8] for the proof of this claim. We observe that in most of the previous works dealing with quaternionic Möbius transformations, with some exceptions as, for example, [3, 6], one considers coefficients on the left which lead to a corresponding left action. We have chosen to use coefficients and corresponding action on the right to better resemble slice regular maps. Left multiplication by a constant quaternion is not slice regular but right multiplication is.
On the other hand, [16] introduced the regular counterpart of linear fractional transformations (see also [12]). For every $A\in\mathrm{Sp}(1,1)$ as above the regular Möbius transformation of $\mathbb{B}$ associated to $A$ is the (regular) map $\mathcal{F}_{A}:\mathbb{B}\rightarrow\mathbb{B}$ given by
@@
Every regular Möbius transformation is a homeomorphism of $\mathbb{B}$. However, as a consequence of the properties of the regular product, the assignment $A\mapsto\mathcal{F}_{A}$ is not a group homomorphism. Nevertheless, these are the transformations that better fit the analysis of slice regular functions. We will denote by $\mathcal{M}(\mathbb{B})$ the set of all regular Möbius transformations. In particular, $\mathcal{M}(\mathbb{B})$ is not a group for the usual composition of functions. However, it is interesting to note that $\mathcal{M}(\mathbb{B})$ is the orbit of the identity map on $\mathbb{B}$ for a suitable $\mathrm{Sp}(1,1)$-action: see [12, Corollary 9.17] for further details. In other words, $\mathcal{M}(\mathbb{B})$ does have a close relationship with a Lie group, namely $\mathrm{Sp}(1,1)$.
Whenever we need to clearly distinguish the transformations of $\mathcal{M}(\mathcal{B})$ with those obtained from the $\mathrm{Sp}(1,1)$-action considered above, we will refer to the latter as non-regular Möbius transformations.
@@
A map $\mathcal{F}:\mathbb{B}\rightarrow\mathbb{B}$ is a regular Möbius transformation if and only if there exist $a\in\mathbb{B}$ and $u\in\mathrm{Sp}(1)$ such that
@@
    $\mathcal{F}_{a}$ is a diffeomorphism whose directional derivative at $a$ in the direction of $\alpha$ is given by
[…]
@@
Let us fix $a\in\mathbb{B}$ and consider the functions on $\mathbb{B}$ defined by the expressions
$f(q)=q\overline{a}-1,\quad g(q)=q-a,$
so that we have
$\mathcal{F}_{a}=\frac{1}{f^{s}}f^{c}*g.$
From the definitions we can compute
$f^{s}(q)$ $=(q\overline{a}-1)*(qa-1)=q^{2}|a|^{2}-2q\operatorname{Re}(a)+1$
$\big{(}f^{c}*g\big{)}(q)$ $=q^{2}a-q(a^{2}+1)+a,$
for every $q\in\mathbb{B}$.
On the other hand, we note that either $f\equiv-1$ (for $a=0$), or we have $f(q)=0$ if and only if $q=(\overline{a})^{-1}$. Since $|(\overline{a})^{-1}|>1$, we conclude that $f$ does not vanish in $\mathbb{B}$, and so [12, Proposition 3.10] implies that $f^{s}$ does not vanish in $\mathbb{B}$ either. Hence we obtain the formula for $\mathcal{F}_{a}$ in the statement and also conclude that its expression is well-defined: i.e. its denominator does not vanish in $\mathbb{B}$.
From the previous remarks, we can evaluate to conclude that $\mathcal{F}_{a}(a)=0$ and $\mathcal{F}_{a}(0)=a$. For $a=0$, it is straightforward to see that $\mathcal{F}_{0}(q)=-q$, for all $q\in\mathbb{B}$. In particular we obtain (1).
For the first claim in (2) we use Corollary 8.24 and Theorem 9.12 from [12]. The latter implies that $\mathcal{F}_{a}$ is a homeomorphism, in particular injective, and so the former implies that $\mathcal{F}_{a}$ is a local diffeomorphism. It follows that $\mathcal{F}_{a}$ is a diffeomorphism.
To obtain the formula in (2) we compute as follows
[…]
$\quad\cdot\big{(}(a+t\alpha)^{2}a-(a+t\alpha)(a^{2}+1)+a\big{)}\bigg{)}$
[…]
$\quad\cdot\big{(}a^{3}-a(a^{2}+1)+a\big{)}$
$+\big{(}a^{2}|a|^{2}-2a\operatorname{Re}(a)+1\big{)}^{-1}\cdot$
[…]
$=$ $\;\big{(}a^{2}|a|^{2}-a^{2}-|a|^{2}+1\big{)}^{-1}(a\alpha a-\alpha)$
$=$ $\;\big{(}(1-a^{2})(1-|a|^{2})\big{)}^{-1}(a\alpha a-\alpha),$
which yields the stated expression. We used the chain rule in the first identity. In the computations that follow, we used Leibniz rule for the product (in $\mathbb{H}$) of $\mathbb{H}$-valued functions, which is justified from the bilinearity of such product. ∎